Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

% complete % complete
units DM conversion work in process, April 1 18,000 38% 71% units started during April 80,000 work in process, April 30 44,000 28% 74%

The cost of beginning work in process and the costs added during April
were as follows:

DM Conversion Total cost work in process, April 1 $ 64,996 $152,785 $217,781 costs added during April 190,336 232,407 422,743 total costs 255,332 385,192 640,524

Calculate the total cost of the handling department's work in process
inventory at April 30 using the weighted average process costing method.
Answer #1

Unit transferred out = 18000+80000-44000 = 54000

Equivalent unit of material = 54000+44000*28% = 66320

equivalent unit of conversion = 54000+44000*74% = 86560

Cost per equivalent unit of material = 255332/66320 = 3.85

Cost per equivalent unit of conversion = 385192/86560 = 4.45

Cost of handling department's work in process
inventory at April 30 using the weighted average process costing method = (12320*3.85+32560*4.45) = 192324